finance


Carlyle Inc. is considering two mutually exclusive projects. Both require an initial investment of $15,000 at t = 0. Project S has an expected life of 2 years withafter-tax cash inflows of $7,000 and $12,000 at the end of Years 1 and 2, respectively. Project L has an expected life of 4 years with after-tax cash inflows of $5,200at the end of each of the next 4 years. Each project has a WACC of 10.75%, and neither can be repeated. The controller prefers Project S, but the CFO prefers ProjectL. How much value will the firm gain or lose if Project L is selected over Project S, i.e., what is the value of NPVL - NPVS?
Answer #1

For Project S find the present value of the cash inflows:

7,000/1.1075 + 12,000/1.1075 = 6320.54 + 9783.49 = 16104.03

Then subtract the initial cost: 16104.03 – 15,000 = 1104.03

So the NPV of Project S is 1,104.03

For Project L, find the present value of the cash flows:

Plug into financial calculator: Number of periods = 4, Interest rate = 10.75, Payment = -5200, Future value =0, end of period (ordinary annuity), and solve for present value and get: 16,219.21

Then subtract the initial cost: 16,219.21 – 15000 = 1,219.21

So the NPV of Project L is 1,219.21.

If they choose L, they will lose 115.18.

NPVL – NPVS = 1,219.21 – 1,104.03 = 115.18

answered by: Alan Burt